# Chat-3D v2: Bridging 3D Scene and Large Language Models with Object Identifiers

### 13 Dec 2023 v1

@@ 2.2 Multi-modal Large Language Models
3D point cloud is a fundamental visual modality. Recently, numerous attempts are made to learn discriminative and robust representations for point cloud objects. PointBERT [39], Point-MAE [25], Transformer-OcCo [30], and Point-m2ae [43] employ self-supervised learning approaches to extract meaningful representations of 3D objects from unlabeled point cloud data. Another series of works [36, 22, 44, 10, 35, 34] aims to extend representation from other modalities to 3D. For instance, ULIP [36] and OpenShape [22] construct (3D-image-text) triplets to align point clouds within the CLIP [27] representation space. These powerful 3D representations can effectively capture the detailed information of a 3D object. We segment the 3D scene into objects and extract features for each object, which yields a set of object features to represent the 3D scene.
@@ 5 Experiments
Figure 3: Example of 3D-language scene understanding tasks.
5.1 Implementation Details

### 15 Dec 2023 v2

@@ 2.2 Multi-modal Large Language Models
Recently, numerous efforts have been made to learn discriminative and robust representations for 3D point clouds, which serve as a fundamental visual modality. Approaches such as PointBERT [39], Point-MAE [25], Transformer-OcCo [30], and Point-m2ae [43] employ self-supervised learning techniques to extract meaningful representations of 3D objects from unlabeled point cloud data. Another set of works [36, 22, 44, 10, 35, 34] seeks to extend representation from other modalities to the 3D domain. For example, ULIP [36] and OpenShape [22] construct 3D-image-text triplets to align point clouds within the CLIP [27] representation space. C-MCR [35] and Ex-MCR [34] learn contrastive representations between various modalities, including 3D point clouds. They leverage knowledge from existing MCR spaces to tackle to challenge of lacking paired data. These robust 3D representations effectively capture detailed information about 3D objects. Our approach involves segmenting the 3D scene at the instance level and extracting a set of object features to represent the entire scene.
@@ 3.3 Scene-level Alignment
Figure 3: Example of 3D-language scene understanding tasks.